# August 2025 IPMU25-0038

### v1

@@ 9 Conclusion
Single-field models of inflation, with or without PBH production, as well as inflation models based on $F(R)$ gravity, are under scrutiny due to CMB observations with increasing precision. On the theoretical side, the single-field models of inflation are also constrained by quantum gravity conjectures known as Swampland program [50], though those constraints do not rule out Starobinsky inflation yet [51, 4].
The recent ACT/DESI observational data disfavors the Starobinsky inflation model (1) by $2\sigma$ and thus motivates a search for its modifications. The simplest modifications including only one additional term proportional to $R^{3}$ or $R^{4}$ can provide the desired increased value of the cosmological tilt $n_{s}$ of scalar perturbations but the running index $\alpha_{s}$ remains negative [18, 19]. To obtain a positive value of $\alpha_{s}$, one should consider further modifications of the Starobinsky model. Generically, it leads to complicated inflaton potentials $V_{E}$ that can be studied only numerically. To overcome this difficulty, our new version of the SR approximation in the Jordan frame allows one to investigate the $F(R)$ gravity models of inflation without resorting to the Einstein frame. Using this SR approximation, we proposed and studied a new inflation model defined by Eq. (113) that agrees with the ACT constraints to the CMB observables for some values of the model parameters. To the best of our knowledge, the proposed model is the first $F(R)$-gravity-based model of inflation with a positive running index $\alpha_{s}$ and $1\sigma$-agreement with the ACT/DESI data. The other inflationary parameters are in agreement with the observations as well.
Our approach is also applicable to other $F(R)$ gravity models of inflation, including those with PBH production [31, 32] and those combining inflation with dark energy [52, 53, 54, 55].
Since the difference between the Planck and ACT observations of CMB is just about $2\sigma$, it is still premature to dispose the Planck data. Therefore, future precision measurements of the cosmological tilts and their running by experiments such as LiteBIRD [56] and the Simons Observatory [57] are needed to resolve the current tensions between the available data.

### v2

@@ 10 Conclusion
Single-field models of inflation, with or without PBH production, as well as inflation models based on $F(R)$ gravity, are under scrutiny due to the CMB observations with increasing precision. On the theoretical side, the single-field models of inflation are also constrained by quantum gravity conjectures known as the Swampland program [60], though those constraints do not rule out the Starobinsky inflation [61, 6].
The $\mu$-type distortion [62, 63] of the CMB leads to additional constraints on the primordial power spectrum at smaller scales up to $k\sim 10^{4}\,\mathrm{Mpc}^{-1}$. The $\mu$-distortion was estimated in Refs. [64, 65, 66, 67, 68]. The upper bound from COBE/FIRAS observations [69] is given by $\mu<9\cdot 10^{-5}$ ($95\%\,\text{C.L.}$). As regards small deformations of the Starobinsky model, it was found in Ref. [70] that $\mu\simeq 10^{-10}$ that is significantly below the observational limit.
The recent ACT/DESI observational data disfavors the Starobinsky inflation model (1.1) by $2\sigma$ and thus motivates a search for its modifications. The simplest modifications including only one additional term proportional to $R^{3}$ or $R^{4}$ can provide the desired increased value of the cosmological tilt $n_{s}$ of scalar perturbations but the running index $\alpha_{s}$ remains negative [18, 19]. To obtain a positive value of $\alpha_{s}$, one should consider further modifications of the Starobinsky model. Generically, it leads to a complicated inflaton potential $V_{E}$ that can be studied only numerically. To overcome this difficulty, our new version of the SR approximation in the Jordan frame allows one to investigate the $F(R)$ gravity models of inflation without resorting to the Einstein frame. Using this SR approximation, we have proposed and studied a new inflation model defined by Eq. (8.3) that agrees with the ACT constraints to the CMB observables for some values of the model parameters. To the best of our knowledge, the proposed model is the first $F(R)$-gravity-based model of inflation with a positive running index $\alpha_{s}$ and $1\sigma$-agreement with the ACT/DESI data. The other inflationary parameters are in agreement with the observations as well.
As was demonstrated in section 9, our approach is extendable to unified $F(R)$ gravity models of inflation, PBHs production and DE, without resorting to the Einstein frame, see Refs. [71, 72, 73] for comparison.
Since the difference between the Planck and ACT observations of the CMB is just about $2\sigma$, it is still premature to dispose the Planck data. Therefore, future precision measurements of the cosmological tilts and their running by experiments such as LiteBIRD [74] and the Simons Observatory [75] are needed to resolve the current tensions between the available data.